# Masked Autoencoders are PDE Learners

### v2

@@ C.2 Additional 2D MAE Predictions
C.3 2D Smoke Buoyancy Predictions
Figure 9: MAE validation reconstructions after training on 2D Navier-Stokes data with variable buoyancy factors [18]. The MAE model is trained on a resolution of $(n_{t},n_{x},n_{y})=(56,128,128)$ with three data channels $(\rho,v_{x},v_{y})$ and a masking ratio of 0.75. Triplets are shown with the masked input (top), MAE reconstruction (middle), and ground truth (bottom), with the top, middle, and bottom triplets displaying density, X velocity, and Y velocity. The complex dynamics is challenging; indeed, many of the fine details are lost in the MAE reconstruction. We train with a larger model (45M params) and patch size $(2,8,8)$, which takes around 9 hours on a NVIDIA RTX A6000 GPU.
@@
Appendix D Training Details
D.1 PDE Feature Prediction
[…]
D.2 Time-stepping
Time-stepping is performed autoregressively by predicting multiple timesteps simultaneously to reduce error accumulation [8]. Furthermore, the pushforward trick [8] is implemented. This adds model noise to inputs during training by making a prediction of a future timestep and using that prediction as a noised input the model; importantly gradients are not calculated for the initial pass. Hyperparameters used during time-stepping are presented in Table 9(c). For each seed, an experiment in 1D takes around 30 minutes, or 45 minutes if fine-tuning an encoder on a NVIDIA RTX 2080 Ti; an experiment in 2D takes around 15 minutes, or 2 hours if fine-tuning an encoder on a NVIDIA RTX A6000.
D.3 Super-resolution
Super-resolution is performed in multiple time steps simultaneously throughout the low-resolution PDE trajectory. We present a detailed schematic of the SR pipeline in Figure 10. Since the encoder operates on low-resolution physics, we use the multi-resolution MAE in 1D. In 2D, the lowest resolution used during multi-resolution training is too large, so a separate MAE is trained on $32\times 32$ data. The discretization inversion operator $D^{-1}$ used is linear interpolation in 1D and bicubic interpolation in 2D. The hyperparameters used for the Resnet encoder, which uses residual dense blocks [66], and FNO operator are reported in Table 9(c). Super-resolution tasks tend to be fast, on the order of minutes for 1D and 2D samples, and are generally limited by dataloading and memory considerations.
@@ E.2 Time-stepping
E.3 Super-resolution

### v3

@@ B.2 Additional 2D MAE Predictions
B.3 2D Smoke Buoyancy Predictions
Figure 7: MAE validation reconstructions after training on 2D Navier-Stokes data with variable buoyancy factors [Gupta and Brandstetter, 2022]. The MAE model is trained on a resolution of $(n_{t},n_{x},n_{y})=(56,128,128)$ with three data channels $(\rho,v_{x},v_{y})$ and a masking ratio of 0.75. Triplets are shown with the masked input (top), MAE reconstruction (middle), and ground truth (bottom), with the top, middle, and bottom triplets displaying density, X velocity, and Y velocity. The complex dynamics is challenging; indeed, many of the fine details are lost in the MAE reconstruction. We train with a larger model (45M params) and patch size $(2,8,8)$, which takes around 9 hours on a NVIDIA RTX A6000 GPU.
@@
Appendix C Training Details
Hyperparameters used for the FNO and Unet models during time-stepping are presented in Table 9(c). Additionally, for the SR pipeline, hyperparameters used for the Resnet encoder, which uses residual dense blocks [Zhang et al., 2018], and FNO operator are reported in Table 9(c). We present a schematic of the SR pipeline in Figure 8.
@@ Appendix C Training Details
Appendix D Latent Arithmetic
MAE encoders have shown strong capabilities in extracting information from self-supervised PDE learning, creating opportunities to operate in this latent space. This could be beneficial since many PDEs are compositions of simpler phenomena, and recombining PDE solutions in latent space may result in obtaining novel PDE solutions for free. After pretraining on the 1D KdV-Burgers equation, we consider an arithmetic task where samples of the Heat and Burgers equation are embedded and added in the latent space before being decoded to a novel solution (Figure 9). Concretely, we generate 1D Heat data from the PDE: $\partial_{t}u-\nu\partial_{xx}u=0$, 1D inviscid Burgers data from the PDE: $\partial_{t}u+u\partial_{x}u=0$, and 1D viscous Burgers data from adding the two PDEs: $\partial_{t}u-\nu\partial_{xx}u+u\partial_{x}u=0$. When given identical initial conditions and coefficients, PDE reconstructions obtained from summing latent Heat and Burgers embeddings qualitatively resemble solutions of the viscous Burgers PDE. In addition, if different latent embeddings can be added, the weighting of each embedding can be varied to control the resemblance of the reconstruction to make interpolated samples that have more shock formation or diffusive behavior (e.g., more resemble the Burgers or Heat equation).
@@ E.2 Time-stepping
E.3 Comparison to Contrastive Learning with Lie Augmentations
Table 15: We compare our approach to a contrastive self-supervised approach. After training a masked and contrastive encoder on the KdV-B pretraining set, we compare conditioning an FNO backbone to time-step different downstream 1D PDEs. The MAE encoder shows comparable performance within the pretraining set (KdV-B), but has better generalization behavior to unseen PDEs. Validation errors are reported as normalized L2 loss summed over all PDE timesteps
Model KdV-B Heat Burgers Adv
FNO 1.506 0.827 1.386 0.567
FNO-Contrastive 1.171 0.918 0.916 0.555
FNO-MAE 1.183 0.721 0.831 0.299
E.4 Super-resolution
@@
Appendix F Comparison of Latent Embeddings
To understand the versatility of masking pretraining, we would like to consider what it means to be a PDE learner beyond predicting future physics. For example, when reasoning about physics, humans can add, subtract, and rearrange equations to derive new knowledge. Additionally, we are able to identify what similar or different physical phenomena are in order to reason about relationships between physical quantities. Lastly, when observing physics, we intuitively understand that solutions must evolve forward with time and be temporally coherent. In the context of machine learning, these behaviors would have to be manifested in latent representations of physical solutions, and to this end, we demonstrate that masked autoencoders can learn expressive, general representations of PDEs.
To do this, we propose a set of experiments to compare the MAE model against other pretraining paradigms such as contrastive or transfer learning. Firstly, we pretrain and MAE or contrastive encoder [Mialon et al., 2023] on the 1D KdV-Burgers pretraining set. Additionally, we pretrain a FNO and Unet model to predict future timesteps on this dataset in order to model a transfer learning scenario. Given these pretrained models, we seek to understand their latent representations. To do this, we embed PDE samples from three different scenarios.
Firstly, we embed equation samples from the Heat and Burgers equations and add these two embeddings in latent space; the average pairwise distance is calculated between this summed embedding and a corresponding embedding from the viscous Burgers equation, which is an experiment in latent arithmetic (Arithmetic). Next, the average pairwise distance between embeddings from the Heat equation is calculated; embedded samples have varying initial conditions but evolve with the same coefficients, which tests if models can embed similar dynamics to similar representations (Similarity). Lastly, he average pairwise distance between embeddings of subsequent timesteps of the Heat equation is calculated, measuring temporal coherence of latent embedding (Temporal). To ensure a fair comparison, embeddings are projected to a common dimension d = 16 with PCA and normalized ($v=v/max(||v||_{2})$). Masked pretraining performs well across these experiments and learns general representations due to its minimal inductive bias, compared to contrastive learning or neural solvers which have specific objectives to either maximize similarity or predict next the timestep.
Table 18: Models are pretrained on the same KdV-B dataset and used to produce latent embeddings across different scenarios. The average pairwise L2 distance is reported in each case.
Model Arithmetic $\downarrow$ Similarity $\downarrow$ Temporal $\downarrow$ Average $\downarrow$
MAE 0.5072 1.3674 0.5277 0.8007
Contrastive 1.4839 1.3131 0.8626 1.2199
FNO 0.9442 1.3562 0.7565 1.0190
Unet 0.9447 1.3562 0.7397 1.0135